PROBLEM 7.82 - 100 MPa Solve Prob. 7.81, using the maximum-shearing-stress criterion. PROBLEM 7.81 The state of plane stress shown occurs in a machine component made of a steel with oy = 325 MPa. Using the maximum-distortion-energy criterion, determine whether yield will occur when (a) o = 200 MPa, (b) o = 240 MPa, (C) 00 = 280 MPa. If yield does not occur, determine the corresponding factor of safety.
Answer #1

200 16 (CL = loo - 1-23 264.57 325. 29S-91 F. o . s = 1 098 dees ocuu6m O tar same materre 2. 2 2- 2 2 ぬ= .as-162.. ç MPa = }.625め 2 2. 2. lSD 2 2. 2